LaylaSanaii:

Write an expression that is equivalent to two over three (4x + 9).

Ember:

Basically, I think the expression you are referring to is, \[\frac{ 2 }{ 3 }(4x+9)\] Do you have any answer choices or can you write anything you'd like?

LaylaSanaii:

yea

LaylaSanaii:

1- eight over threex + 6 2- eight over threex + 9 3- two over threex + 6 4- two over threex + 9

LaylaSanaii:

those are the answer choices i think 2 or 4

LaylaSanaii:

thx

Adventuretime:

you would multiply 2 and three and the 4 then add nine hopfully this helps

Vocaloid:

Distributive property A(B+C) = AB + AC In other words, take the number outside the parentheses, multiply it by the first number Repeat with the second number Add the results